7) Given f(x,y)= x^2+y^2+2, subject to the constraint g(x,y)=x^2+xy+y^2-4=0, write the system of equations which must be solved to optimize f using Lagrange Multipliers.
Answer #1

- Given f(x,y) = x² tý tz subject to the construint g(x,y) = x² + xy + y - 40 0 Construct a Lagrange function 162,5) = f(1,3)P. 2 ::(+2y) = y(x+y) x trung= 2ang ty? v = ... = . - Piut y = x in D, + x +7 _4=; 7:3x? -4 = 0 -34-4 are are critical point